OpenStudy (anonymous):

The length of a rectangular label is 5 cm more than the width. The perimeter is 50 cm. Find the exact width and length

OpenStudy (anonymous):

2l + 2w = 50 l = w + 5 so 2(w+5) + 2w = 50 4w = 40 w = 10 so l = 15 length = 15 cm width = 10 cm